OpenStudy (anonymous):

Solve the equation. Show your work. 8x – 12 = 4x + 24 8x – 12 = 4x + 24 -4 -4x 4x -12 =+ 24 +12 12 4x 36 4x/4 36/4= 9

OpenStudy (misssunshinexxoxo):

Simplifying 8x + -12 = 4x + 24 Reorder the terms: -12 + 8x = 4x + 24 Reorder the terms: -12 + 8x = 24 + 4x Solving -12 + 8x = 24 + 4x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-4x' to each side of the equation. -12 + 8x + -4x = 24 + 4x + -4x Combine like terms: 8x + -4x = 4x -12 + 4x = 24 + 4x + -4x Combine like terms: 4x + -4x = 0 -12 + 4x = 24 + 0 -12 + 4x = 24 Add '12' to each side of the equation. -12 + 12 + 4x = 24 + 12 Combine like terms: -12 + 12 = 0 0 + 4x = 24 + 12 4x = 24 + 12 Combine like terms: 24 + 12 = 36 4x = 36 Divide each side by '4'. x = 9 Simplifying x = 9 You got it :)

OpenStudy (anonymous):

Did i write it right?

OpenStudy (anonymous):

Yeah i got what you wrote just not sure if mines is wrong or right.

OpenStudy (anonymous):

Because yours is a lot longer seems like i missed something XD

OpenStudy (misssunshinexxoxo):

It is correct

OpenStudy (anonymous):

Alright thanks Missunshine